OpenStudy (anonymous):

Use the quadratic formula to find the x-intercepts of the function. Round to the nearest tenth if necessary. y = -2x2 + 5x - 1

OpenStudy (aaronandyson):

"The x-intercepts are the points where the curve crosses the x-axis and the y-coordinates are zero. Let y = 0, then you get 2x2+5x-1=0 Use the quadratic formula with a = -2, b = 5, and c = -1 to find x."

OpenStudy (anonymous):

A. {2.3, 0.2} B. {2.2, -5.8} C. {-9.2, -0.8} D. no real solution

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

\[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a}\]